Lori decides to paint Marie's room a shede of blue. She measures the height and width of each wall in the rectangular room. She finds that two walls have a width of 12 feet and the other two have a width of 14 feet. The ceiling is 9 feet high. Find the area of each wall. Find the total area of all four walls plus the ceiling.
Answer by gonzo(654)

height of each wall is 9 feet.
area of 14 foot wide wall = 14 * 9 = 126 sqft.
area of 12 foot wide wall = 12 * 9 = 108 sqft.
total area of walls is 2 * (126 + 108) = 2 * 234 = 468 sqft.
area of ceiling is 14 * 12 = 168 sqft.
sqft is square feet.
total area of 4 walls plus ceiling is 468 + 168 sqft = 636 sqft.